OpenStudy (anonymous):

find the special product:(2y-3)^2

OpenStudy (anonymous):

If you managed the other one, this is easy in comparison...:-)

OpenStudy (anonymous):

is it like the foil ones?

OpenStudy (anonymous):

(2y-3)(2y-3) = 2y(2y-3) - 3(2y-3) (Except you ought to be able to just do it in your head)

OpenStudy (anonymous):

i just get confused when it says special product o.o

OpenStudy (anonymous):

Me too, I have no idea what they are talking about......

OpenStudy (anonymous):

The only thing is (a-b)^2 = a^2 -2ab +b^2 is worth just knowing

OpenStudy (anonymous):

If you do enough of them you will know anyway.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

K